the vertex angle of an isosceles (a+15) one of the base angles measures 7a. Find a and each angle measure
Answer by jim_thompson5910(35256)


(a+15) + (7a) + (7a) = 180

a+15 + 7a + 7a = 180

15a+15 = 180

15a = 180-15

15a = 165

a = 165/15

a = 11

a+15 = 11+15 = 26

7a = 7(11) = 77

So in the end,

a = 11

a+15 = 26

7a = 77

Therefore, the three angles are 26, 77, and 77 degrees.